Max points offered for a solution- THANKS!
Answer #1

apply dsintheta = (m+0.5)L

for highest order theta = 90 deg

so m +0.5= d/L

m = 1500nm/690 nm

m = 2.17-0.5 = 2

highest order = 2
Answer #2

sin(theta) = m*lambda/a

sin(theta) = 1 (max)

m=1*a/lambda = 1500/690 = 2.17

m = 2 (2nd order dark fringe)
Answer #3

dsin(theta)=m*lamda

sin(theta)<=1

m*lamda/d <=1

m<=2.19

m= 2